OpenStudy (anonymous):

If f(x) = x + 2, find f(3).

OpenStudy (nubeer):

replace x with 3

OpenStudy (anonymous):

f(3) = 3+2 > f(3)= 5

OpenStudy (anonymous):

f(3)=x+2

OpenStudy (anonymous):

yup!

OpenStudy (nubeer):

yes and the x in the equation too.

OpenStudy (anonymous):

Thanks @InstagramDoll